the population is normally distri a) (9.22, 9.78] b) 19.24, 9.76) (9.28, 9.72] 19.30, 9.70 8 is too high for his district, so he randomly selects 50 injuries and finds that 17 occurred during practices. 25) A report indicates that 60% of football injuries occur during practices. A head trainer daims that this What test would you use test his claim? a) z-test for mean b) t-test for mean c) z-test for proportion d) none of these
Answer #1

Given, n=50, p=60/100=0.60 and X=17.

Here, np=50*0.60=30 and n(1-p)=50(1-0.60)=20 are more than or equal to 10. Hence, we should use the z-test for proportion. S, the correct answer is "C".